OpenStudy (phebe):

Can the set of lenghs be side of a right angle? ~16cm,63cm,66cm A-yes B-no

OpenStudy (campbell_st):

use pythagoras theorem to check does \[66^2 = 63^2 + 16^2\]

OpenStudy (phebe):

so do i subtract

OpenStudy (anonymous):

Just see if that equation is correct. If it is, then the set of lengths can be sides of a right angle. If the equation is not correct, then the set of lengths cannot be sides of a right triangle

OpenStudy (phebe):

um... do u know whether its yes or no

OpenStudy (campbell_st):

no... square all terms... then see if the sum of the right hand side is the same as the left hand side

OpenStudy (phebe):

ok

OpenStudy (campbell_st):

the question is basically asking does 66^2 have the same value as 63^2 + 16^2

OpenStudy (phebe):

i know waht the question is asking so is it yes or no

OpenStudy (campbell_st):

well if you calculate it... you'll find out...

OpenStudy (campbell_st):

or you can wait and someone always gives the answer

OpenStudy (phebe):

:( but im in a hurry tho

OpenStudy (campbell_st):

here is an online calculator... it will speed things up http://web2.0calc.com/

OpenStudy (anonymous):

nirma will give u the ansewer now

OpenStudy (phebe):

ik yayay

OpenStudy (anonymous):

hurry nirma

OpenStudy (nirmalnema):

for right triangle (hypo)^2=(base)^2+(perpen)^2 hypotenuse is the longest side.. 66^2=63^2 + 16^2 4365 is not equal to 4225 so it is not right angle tri.

OpenStudy (campbell_st):

there you go... the answer...

OpenStudy (phebe):

the answer is no rit

OpenStudy (campbell_st):

that's correct

OpenStudy (nirmalnema):

@kx2bay y u r in so hurry its not ur que ??

OpenStudy (phebe):

um... wats que

OpenStudy (anonymous):

yeah wats que

OpenStudy (phebe):

ikr lol